multi-step equations with fractions and decimals.
1.- 0.7n - 1.5 + 7.3n= 14.5

2.- 16.3k + 19.2 + 7.5k= -64.1

Answer by malakumar_kos@yahoo.com(315)


multi-step equations with fractions and decimals.
1.- 0.7n - 1.5 + 7.3n= 14.5
2.- 16.3k + 19.2 + 7.5k= -64.1

solution:
1) -0.7n-1.5+7.3n = 14.5
n(7.3-0.7) = 14.5+1.5
n(6.6) = 16.0
n = 16/6.6 = 160/66 = 80/33

2) -16.3k+19.2+7.5k = -64.1
k(-16.3+7.5) = -64.1-19.2
k(-8.8) = -83. 3
k = 83.3/8.8 = 833/88